OpenStudy (goformit100):

Check Please is my answer correct ? If A = {1, 2, 3} and f , g are relations corresponding to the subset of A × A indicated against them, which of f , g is a function? Why? f = {(1, 3), (2, 3), (3, 2)} g = {(1, 2), (1, 3), (3, 1)

OpenStudy (anonymous):

use substitution to solve system of equations? y=x+5 3x+y=25

OpenStudy (anonymous):

I NEED HELP BRUH

OpenStudy (goformit100):

I got f is a function since each element of A in the first place in the ordered pairs is related to only one element of A in the second place while g is not a function because 1 is related to more than one element of A, namely, 2 and 3 is it correct.

OpenStudy (anonymous):

HOW WOULD YOU DO THAT SHOWING WORK THOUGH.

OpenStudy (anonymous):

use substitution to solve system of equations. y=x+5 3x+y=25